OpenStudy (avengedslipknot):

what is the area of a rectangular octagon with an apothem 16 inches long and a side 19 inches long. Round to the nearest inch. (PLS HELP)

OpenStudy (michele_laino):

here, we have to apply this formula: \[area = \frac{{perimeter \times apothem}}{2}\]

OpenStudy (michele_laino):

wherein: \[perimeter = 19 \times 8 = ...?\]

OpenStudy (avengedslipknot):

456

OpenStudy (king_of_kings):

ur right